# Barbie: Text to Barbie-Style 3D Avatars

### v2

@@ 3.4 Apparel Generation
Apparel Initialization. With the body topology retained by our proposed self-evolving human prior loss, we can easily provide semantically aligned initialization for various apparel. Specifically, we utilize the human body segments of SMPL-X and pre-defined masked templates to cover more than a dozen types of daily clothing and accessories. By cropping the corresponding sub-mesh according to the apparel type, we obtain a closed template mesh $M_{\text{temp}}$ to initialize the apparel DMTet through geometry processing such as scaling and sewing. For further details, please refer to the supplementary material.
Apparel Geometry Modeling. Similar to human body geometry generation, we utilize human-specific diffusion models to optimize outfits geometry with the following SDS loss:
@@
$\nabla_{{\theta}_{\text{a}}}\mathcal{L}^{\text{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hd}}}(d^{\text{a+h}}_{t};y_{\text{a+h}},t)-\epsilon)\frac{\partial d^{\text{a+h}}}{\partial{\theta}_{\text{a}}}\right],$ (8)
where ${\theta}_{\text{a}}$ is the apparel DMTet parameters, $y_{\text{a+h}}$ is the text description of the appareled avatar, and $n^{\text{a+h}}$ and $d^{\text{a+h}}$ are the rendered appareled human normal and depth maps, respectively. However, human-specific diffusion models are not well-suited for creating clothes and accessories, as they lack of in-domain details and diversity of generated outfits. Hence, we additionally introduce the object-specific diffusion models proposed by RichDreamer (Qiu et al. 2024), which are trained on large-scale LAION dataset (Schuhmann et al. 2022).
@@
where $y_{\text{a}}$ is the text description of a single garment or accessory, and $n^{\text{a}}$ and $d^{\text{a}}$ are the rendered normal and depth maps of the apparel in the undressed state. In this way, the object-specific diffusion models provide a powerful geometry prior for generating high-fidelity clothes and accessories.
Template-Preserving Loss. Similar to human-specific diffusion models, relying on the object-specific diffusion models may lead to geometric artifacts such as holes (Sec. 4.4). To address this problem, we propose a template-preserving loss that enforces the apparel geometry to cover the template mesh $M_{\text{temp}}$ during the optimization process to ensure geometric integrity. The loss function is formulated as follows:
$\mathcal{L}_{\text{temp}}=\sum_{p\in P_{\text{temp}}}max(0,s_{\theta_{\text{a}}}(p)),$ (10)
@@
Specifically, this loss guarantees that the apparel geometry covers the template mesh by pushing the sampled SDF values of $s_{\theta_{\text{a}}}$ inside the template mesh to be negative. Instead of employing the MSE function to provide prior constraints such as Eq. (4), the template-preserving loss only restricts the SDF values of the points inside the template, ensuring geometric rationality without limiting the sculpting of geometric details outside the template mesh.
@@
$\mathcal{L}_{\text{app-geo}}=\mathcal{L}^{\text{hn}}_{SDS}+\mathcal{L}^{\text{hd}}_{SDS}+\mathcal{L}^{\text{ond}}_{SDS}+\mathcal{L}^{\text{sd}}_{SDS}+\lambda_{\text{temp}}\mathcal{L}_{\text{temp}},$ (11)
where $\mathcal{L}^{\text{sd}}_{SDS}$ is the vanilla Stable Diffusion (SD) SDS loss enforced on the rendered apparel normal maps. As shown in RichDreamer, native SD helps object-specific diffusion models produce more stable results.
@@
$\nabla_{{\theta}_{\text{a}}}\mathcal{L}^{\text{oc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{oc}}}(c^{\text{a}}_{t};d^{\text{a}},y_{\text{a}},t)-\epsilon)\frac{\partial c^{\text{a}}}{\partial{\theta}_{\text{a}}}\right],$ (12)
where $c^{\text{a}}$ is apparel color image. Additionally, human-specific diffusion models are combined to optimize the appearance of outfits in the dressed state to ensure basic texture harmony, and the SDS loss is formulated as:
$\nabla_{{\theta}_{\text{a}}}\mathcal{L}^{\text{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{a+h}}_{t};n^{\text{a+h}},y_{\text{a+h}},t)-\epsilon)\frac{\partial c^{\text{a+h}}}{\partial{\theta}_{\text{a}}}\right].$ (13)
@@
$\mathcal{L}_{\text{app-tex}}=\mathcal{L}^{\text{hc}}_{SDS}+\mathcal{L}^{\text{oc}}_{SDS}+\mathcal{L}^{\text{sd}}_{SDS}.$ (14)
Figure 4: Diverse Range of Barbie-Style Avatar Generation. Rendering color images and normal images for visualization. Please zoom in to see the details. See supplementary materials for video results.
@@ 3.5 Unified Texture Refinement
To address this challenge, we propose a unified texture refinement (UTR) strategy, which jointly fine-tunes the appearance of the assembled avatar under the multi-step SDS (MSDS) proposed by HumanNorm (Huang et al. 2024a). The loss function for UTR is formulated as follows:
$\nabla_{{\theta}_{\text{A}}}\mathcal{L}^{\text{hc}}_{MSDS}=\mathbb{E}_{t,\epsilon}\left[(H(c^{\text{A+h}}_{t};n^{\text{A+h}},y_{\text{A+h}},t)-\epsilon)\frac{\partial c^{\text{A+h}}}{\partial{\theta}_{\text{A}}}\right],$ (15)
where $H$ denotes the multi-step operation function (see HumanNorm for more details), ${\theta}_{\text{A}}=\{\theta_{\text{a}_{i}},i\in[1,...,N]\}$ refers to the parameters of all apparel DMTets, $y_{\text{A+h}}$ is the full text description of the avatar wearing all outfits, and $n^{\text{A+h}}$ and $c^{\text{A+h}}$ are the rendered normal and color images of the avatar wearing all clothes and accessories. Unlike Eq. (14), we only use the human-specific normal-conditioned diffusion model to ensure overall texture style unity. Besides, we use the MSDS loss instead of the native SDS loss to further enhance texture realism, leading to more natural textures. In this way, UTR enhances the texture harmony and consistency of the generated avatar.
@@ 3.6 Implementation Details
Our algorithm is implemented using PyTorch (Paszke et al. 2019) and ThreeStudio (Guo et al. 2023). All experiments are conducted on an Ubuntu server equipped with A6000 GPUs. Generating a human body takes roughly 3 hours with 24GB memory, and generating a cloth or accessory takes roughly 4 hours with 40GB memory. Given a text describing an appareled human, e.g., ”A man wearing X1 and X2”, we first generate a base human body based on ”A man in his underwear”. We then generate the apparel X1 based on ”A piece of X1” and ”A man wearing X1”, and then generate the apparel X2 based on ”A piece of X2” and ”A man wearing X1 and X2”. Finally, fine-tune the overall texture based on the full input text ”A man wearing X1 and X2”. Additional details can be found in supplementary materials.
@@ 4.1 Experimental Settings
Due to space limitations, we present only the most important experimental settings and results here. For complete results and details, please refer to our supplementary materials.
Method B-VQA $\uparrow$ B2-VQA $\uparrow$ GQP $\uparrow$ TAP $\uparrow$
@@
Table 1: Quantitative comparisons with the SOTA methods. The best and second-best results are highlighted in bolded and underlined, respectively. B-VQA: BLIP-VQA. B2-VQA: BLIP2-VQA. GQP: generation quality preference (%), TAP: text-image alignment preference (%).
Baselines. We compare Barbie with the state-of-the-art (SOTA) methods for dressed avatar generation and apparel generation. The compared text-to-avatar methods include DreamWaltz (Huang et al. 2024b), TADA (Liao et al. 2024), X-Oscar (Ma et al. 2024), HumanGaussian (Liu et al. 2024), and HumanNorm (Huang et al. 2024a). The compared text-to-object methods include DreamFusion (Poole et al. 2022), Fantasia3D (Chen et al. 2023), MVDream (Shi et al. 2023), GaussianDreamer (Yi et al. 2024), and RichDreamer (Qiu et al. 2024). Additionally, we provide a qualitative comparison with the latest generation work that has not yet published code in the supplementary material.
Dataset Construction. We utilized ChatGPT to randomly generate 30 text descriptions of appareled avatars, with each example wearing a top, a bottom, a pair of shoes, and two random accessories. 30 descriptions for dressed humans are used to evaluate avatar generation, and 30$\times$5 apparel descriptions are used to evaluate apparel generation. The complete descriptions are listed in the Appendix.
Evaluation Metrics. Existing text-to-3D approaches utilize CLIP-based metrics to evaluate text-image alignment and compare the quality of generated results. However, CLIP-based metrics have been shown (Huang et al. 2023a; Lu et al. 2024) to be insufficient to accurately measure the fine-grained correspondence between 3D content and input prompts, which is further confirmed by experiments in the supplementary material.
Consequently, inspired by Progressive3D (Cheng et al. 2023), we adopt fine-grained text-to-image evaluation metrics including BLIP-VQA (Li et al. 2022, 2023a) and BLIP2-VQA (Li et al. 2023b, a) to evaluate the generation capacity of current methods and Barbie. Specifically, we first convert the prompt into multiple separate questions to retrieve corresponding content, then feed the rendered image of the 3D generated content into the VQA model and ask questions one by one, and finally use the probability of answering “yes” as the evaluation metric. For instance, the input avatar prompt “A man wearing X1.” is converted into “Is the person in the picture a man?” and “Is the person in the picture wearing X1?”. The input apparel prompt “A pair of X1.” is converted into “Is the object in the picture a pair of X1?”. Finally, we randomly select 10 examples from the generated results to conduct a user study and ask 30 volunteers to assess (1) generation quality and (2) text-image alignment, and select the preferred methods.
Figure 5: Qualitative comparisons with the SOTA methods.
@@ 4.2 Comparisons of Dressed Avatar Generation
As shown in the upper part of Tab. 1, our method significantly outperforms the comparison methods across all metrics. The qualitative comparisons illustrated in the upper part of Fig. 5 further highlight the strength of our approach. Compared to other methods, the digital humans generated by Barbie exhibit finer and plausible geometric details, and ensure higher alignment with the input text without omitting any clothing or accessories. Additionally, unlike other compared methods, Barbie’s generation process is flexible and controllable, and the generated bodies, clothes, and accessories can be freely combined.
@@ 4.3 Comparisons of Apparel Generation
As shown in the lower part of Tab. 1, our approach outperforms other methods on most evaluation metrics and is only slightly lower than RichDreamer (the SOTA open-source text-to-3D method) on BLIP-VQA, ranking the second. The qualitative comparisons illustrated in the lower part of Fig. 5 visualize some generation results. Compared to other methods, the clothes and accessories generated by Barbie exhibit more high-fidelity geometry and textures while maintaining high alignment with the text without incorporating incorrect content such as parts of the human body.
Figure 6: Ablation study on (a) self-evolving human prior loss, (b) template-preserving loss, (c) object-specific diffusion models supervision, and (d) unified texture refinement.
@@ 4.4 Ablation Study
Effect of Self-Evolving Human Prior Loss. As shown in Fig. 6-(a), omitting $\mathcal{L}_{\text{prior}}$ results in exaggerated human proportions, greatly reducing the realism of the generated results. Non-evolving human prior loss ensures the rationality of the generated human body, but it leads to over-smooth geometry or limited details. Using the full ${L}_{\text{prior}}$ achieves both detailed and reasonable human body geometry generation.
Effect of Template-Preserving Loss. As shown in Fig. 6-(b), the template-preserving loss is crucial for ensuring the integrity of dress geometry. Omitting $\mathcal{L}_{\text{temp}}$ can result in holes or other geometry artifacts, significantly reducing the quality of the generated apparel.
Effect of Object-Specific Diffusion Models. Fig. 6-(c) shows that introducing an object-specific diffusion model provides powerful guidance for creating finer geometric details and more lifelike appearance.

### v5

@@ 3.4 Apparel Generation
Apparel Initialization. With the body topology retained by our proposed SMPLX-evolving prior loss, we can easily provide semantically aligned initialization for various apparel. Specifically, we utilize the human body segments of SMPL-X and pre-defined masked templates to cover more than a dozen types of daily clothing and accessories. By cropping the corresponding sub-mesh according to the apparel type, we obtain a closed template mesh $M_{\text{temp}}$ to initialize the apparel DMTet through geometry processing such as scaling and sewing. For further details, please refer to the Supp. Mat.
Apparel Geometry Modeling. Similar to human body geometry generation, we utilize human-specific diffusion models to optimize outfits geometry with the SDS loss:
@@
$\nabla_{{\theta}_{\text{a}}}\mathcal{L}^{\text{hd}}_{SDS}$ $=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hd}}}(d^{\text{a+h}}_{t};y_{\text{a+h}},t)-\epsilon)\frac{\partial d^{\text{a+h}}}{\partial{\theta}_{\text{a}}}\right],\vspace{-0.1cm}$ (9)
where ${\theta}_{\text{a}}$ is the apparel DMTet parameters, $y_{\text{a+h}}$ is the text description of the appareled avatar, and $n^{\text{a+h}}$ and $d^{\text{a+h}}$ are the rendered appareled human normal and depth maps, respectively. However, human-specific diffusion models are not well-suited for creating clothes and accessories, as they lack of in-domain details and diversity of generated outfits (Sec. 4.4). Hence, we also introduce the object-specific diffusion models proposed by RichDreamer [59], which are trained on large-scale LAION dataset [67].
@@
$\nabla_{{\theta}_{\text{a}}}\mathcal{L}^{\text{sdn}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{sd}}}(n^{\text{a}}_{t};y_{\text{a}},t)-\epsilon)\frac{\partial n^{\text{a}}}{\partial{\theta}_{\text{a}}}\right],$ (11)
where $y_{\text{a}}$ is the text description of a single garment or accessory, and $n^{\text{a}}$ and $d^{\text{a}}$ are the rendered normal and depth maps of the apparel in the undressed state. Besides, $\mathcal{L}^{\text{sdn}}_{SDS}$ is the vanilla Stable Diffusion (SD) SDS loss enforced on the rendered apparel normal maps. As shown in RichDreamer, native SD helps object-specific diffusion models produce more stable results. In this way, the object-specific diffusion models provide a powerful geometry prior to generating high-fidelity clothes and accessories.
Figure 4: Diverse Range of Barbie-Style Avatar Generation. Rendering color images and normal images for visualization. Please zoom in to see the details and see Supp. Mat. for video results.
Template-Preserving Loss. Similar to human-specific diffusion models, relying on the object-specific diffusion models may lead to geometric artifacts (e.g., holes, see Sec. 4.4). To address this problem, we propose a template-preserving loss that enforces the apparel geometry to cover the template mesh $M_{\text{temp}}$ during the optimization process to ensure geometric integrity. This loss is formulated as follows:
$\mathcal{L}_{\text{temp}}=\sum_{p\in P_{\text{temp}}}max(0,s_{\theta_{\text{a}}}(p)),\vspace{-0.1cm}$ (12)
@@
Specifically, this loss guarantees that the apparel geometry covers the template mesh by pushing the sampled SDF values of $s_{\theta_{\text{a}}}$ inside the template mesh to be negative. Instead of employing the MSE function to provide prior constraints such as Eq. 4, the template-preserving loss only restricts the SDF values of the points inside the template, ensuring geometric rationality without limiting the sculpting of geometric details outside the template mesh.
@@
$\mathcal{L}_{\text{app-geo}}=\mathcal{L}^{\text{hn}}_{SDS}+\mathcal{L}^{\text{hd}}_{SDS}+\mathcal{L}^{\text{ond}}_{SDS}+\mathcal{L}^{\text{sdn}}_{SDS}+\lambda_{\text{temp}}\mathcal{L}_{\text{temp}}.\vspace{-0.1cm}$ (13)
@@
$\nabla_{{\psi}_{\text{a}}}\mathcal{L}^{\text{oc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{oc}}}(c^{\text{a}}_{t};d^{\text{a}},y_{\text{a}},t)-\epsilon)\frac{\partial c^{\text{a}}}{\partial{\psi}_{\text{a}}}\right],\vspace{-0.1cm}$ (14)
$\nabla_{{\psi}_{\text{a}}}\mathcal{L}^{\text{sdc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{sd}}}(c^{\text{a}}_{t};y_{\text{a}},t)-\epsilon)\frac{\partial c^{\text{a}}}{\partial{\psi}_{\text{a}}}\right],\vspace{-0.1cm}$ (15)
where ${\psi}_{\text{a}}$ and $c^{\text{a}}$ is apparel texture field and color image, respectively. Additionally, human-specific diffusion models are combined to optimize the appearance of outfits in the dressed state to ensure basic texture harmony, and the SDS loss is formulated as:
$\nabla_{{\psi}_{\text{a}}}\mathcal{L}^{\text{hc}}_{SDS}=\mathbb{E}_{t,\epsilon}\left[({\epsilon}_{{\phi}_{\text{hc}}}(c^{\text{a+h}}_{t};n^{\text{a+h}},y_{\text{a+h}},t)-\epsilon)\frac{\partial c^{\text{a+h}}}{\partial{\psi}_{\text{a}}}\right].\vspace{-0.1cm}$ (16)
@@
$\mathcal{L}_{\text{app-tex}}=\mathcal{L}^{\text{hc}}_{SDS}+\mathcal{L}^{\text{oc}}_{SDS}+\mathcal{L}^{\text{sdc}}_{SDS}.$ (17)
@@ 3.5 Unified Texture Refinement
To address this challenge, we propose a unified texture refinement (UTR) strategy, which jointly fine-tunes the appearance of the assembled avatar under the following multi-step SDS (MSDS):
$\nabla_{{\psi}_{\text{A}}}\mathcal{L}^{\text{hc}}_{MSDS}=\mathbb{E}_{t,\epsilon}\left[(H(c^{\text{A+h}}_{t};n^{\text{A+h}},y_{\text{A+h}},t)-\epsilon)\frac{\partial c^{\text{A+h}}}{\partial{\psi}_{\text{A}}}\right],\vspace{-0.1cm}$ (18)
where $H(\cdot)$ denotes the multi-step operation function (see HumanNorm for more details), ${\psi}_{\text{A}}=\{\psi_{\text{a}_{i}},i\in[1,...,N]\}$ refers to the parameters of all apparel texture fields, $y_{\text{A+h}}$ is the full text description of the avatar wearing all outfits, and $n^{\text{A+h}}$ and $c^{\text{A+h}}$ are the rendered normal and color images of the avatar wearing all clothes and accessories. Unlike Eq. 17, we only use the human-specific normal-conditioned diffusion model to ensure overall texture style unity. Besides, as the vanilla SDS loss suffers from color oversaturation, we use the MSDS loss to further enhance texture realism, leading to more natural textures. In this way, UTR enhances the texture harmony and consistency of the generated avatar.
@@ 3.6 Implementation Details
Our algorithm is implemented using PyTorch [56] and ThreeStudio [18]. Given a text describing an appareled human, e.g., “A man wearing X1 and X2”, we first generate a base human body based on “A man in his underwear”. We then generate the apparel X1 based on “A piece of X1” and “A man wearing X1”, and then generate the apparel X2 based on “A piece of X2” and “A man wearing X1 and X2”. Finally, fine-tune the overall texture based on the full input text “A man wearing X1 and X2”. Additional details can be found in Supp. Mat.
@@ 4 Experiments
Some examples of 3D body, garments, and accessories generated by Barbie are shown in Fig. 4. Due to space limitations, only the key experimental settings and results are presented here. For complete results and details, please refer to our Supp. Mat.
@@ 4.1 Experimental Settings
Figure 5: Qualitative comparisons with the SOTA text-to-avatar and text-to-3D methods.
Baselines. We quantitatively and qualitatively compare Barbie with the state-of-the-art (SOTA) methods for dressed avatar generation and apparel generation. The compared text-to-avatar methods include DreamWaltz [27], TADA [42], X-Oscar [49], HumanGaussian [44], and HumanNorm [26]. The compared text-to-object methods include DreamFusion [58], Fantasia3D [4], MVDream [69], GaussianDreamer [77], and RichDreamer [59]. Additionally, we conduct a qualitative comparison with recent text-to-disentangled-avatar works [14, 16, 72], though these have not yet released their complete code.
Dataset Construction. We utilized ChatGPT [55] to randomly generate 30 text descriptions of appareled avatars, with each example wearing a top, a bottom, a pair of shoes, and two random accessories. 30 descriptions for dressed humans are used to evaluate avatar generation, and 30$\times$5 apparel descriptions are used to evaluate apparel generation (see the Supp. Mat. for the complete descriptions).
Evaluation Metrics. Existing text-to-3D approaches utilize CLIP-based metrics to evaluate text-image alignment and compare the quality of generated results. However, CLIP-based metrics have been shown [24, 47] to be insufficient to accurately measure the fine-grained correspondence between 3D content and input prompts, which is further confirmed by experiments in the Supp. Mat.
Consequently, inspired by Progressive3D [6], we adopt fine-grained text-to-image evaluation metrics including BLIP-VQA [36, 35] and BLIP2-VQA [37, 35] to evaluate the generation capacity of current methods and Barbie. Specifically, we first convert the prompt into multiple separate questions to retrieve corresponding content, then feed the rendered image of the generated content into the VQA model and ask questions one by one, and finally use the probability of answering “yes” as the evaluation metric. For instance, the input avatar prompt “A man wearing X1.” is converted into “Is the person in the picture a man?” and “Is the person in the picture wearing X1?”. The input apparel prompt “A pair of X1.” is converted into “Is the object in the picture a pair of X1?”. Besides, we randomly select 10 examples from the generated results to conduct a user study and ask 30 volunteers to assess (1) generation quality and (2) text-image alignment, and select the preferred methods.
Method BLIP-VQA $\uparrow$ BLIP2-VQA $\uparrow$ GQP $\uparrow$ TAP $\uparrow$
@@
Table 1: Quantitative comparisons with the SOTA methods. The best and second-best results are highlighted in bolded and underlined, respectively. GQP: generation quality preference (%), TAP: text-image alignment preference (%).
@@ 4.2 Comparisons of Dressed Avatar Generation
Comparison with Text-to-Avatar Methods. As shown in the upper part of Tab. 1, our method significantly outperforms the comparison methods across all metrics. The qualitative comparisons illustrated in the left part of Fig. 5 further highlight the strength of our approach. Compared to other methods, the digital humans generated by Barbie exhibit finer and plausible geometric details, and ensure higher alignment with the input text without omitting any clothing or accessories. Additionally, unlike other compared methods, Barbie’s generation process is flexible and controllable, and the generated bodies, clothes, and accessories can be freely combined.
Figure 6: Qualitative comparisons with the text-to-disentangled-avatar work has not yet released the complete code, and their results are copied from their papers. These methods cannot generate diverse accessories shown in Fig. 4 like our Barbie.
Comparison with Text-to-Decoupled-Avatar Methods. As shown in Fig. 6, our method surpasses TELA [14], LAGA [16], and SO-SMPL [72] in generating refined geometry and lifelike textures. Additionally, Barbie can generate various accessories, including necklaces, glasses, hats, and watches, which are not supported by these methods.
@@ 4.3 Comparisons of Apparel Generation
As shown in the lower part of Tab. 1, our approach outperforms other methods on most evaluation metrics and is only slightly lower than RichDreamer (the SOTA open-source text-to-3D method) on BLIP-VQA, ranking the second. The qualitative comparisons illustrated in the right part of Fig. 5 visualize some generation results. Compared to other methods, the clothes and accessories generated by Barbie exhibit more high-fidelity geometry and textures while maintaining high alignment with the text without incorporating incorrect content, such as parts of the human body.
@@ 4.4 Ablation Study
Effect of SMPLX-Evolving Prior Loss. As shown in Fig. 7-(a), omitting $\mathcal{L}_{\text{prior}}$ results in exaggerated human proportions, greatly reducing the realism of the generated results. Non-evolving human prior loss ensures the rationality of the generated human body, but it leads to over-smooth geometry or limited details. Using the full ${L}_{\text{prior}}$ achieves both detailed and reasonable human body geometry generation.
Effect of Template-Preserving Loss. As shown in Fig. 7-(b), the template-preserving loss is crucial to ensure the integrity of dress geometry. Omitting $\mathcal{L}_{\text{temp}}$ can result in holes or other geometry artifacts, significantly reducing the quality of the generated apparel.
Effect of Object-Specific Diffusion Models. Fig. 7-(c) shows that introducing an object-specific diffusion model provides powerful guidance for creating finer geometric details and a more lifelike appearance.
@@
Figure 7: Ablation study on (a) SMPLX-evolving prior loss, (b) template-preserving loss, (c) object-specific diffusion models supervision, and (d) unified texture refinement.